# An Efficient Image Denoising Method Integrating Multi-resolution Local Clustering and Adaptive Smoothing

### v2

@@ 2.3 Selection of procedure parameters
For our proposed algorithm, we need four parameters: the bandwidth and cutoff for the edge detection step, and bandwidth parameters for the local clustering step and smoothing step. In our implementation, we introduced only one new parameter, the maximum possible axis length of the elliptical neighborhood. From numerical experience, we set this parameter to $6$ pixels for images having resolution less than $100\times 100$, and $10$ pixels for images with higher resolution. Similarly the bandwidth for the clustering method was chosen to be the same as the minimum distance $\gamma$ to apply the method, which we choose to be $3$ pixels for images with resolution less than $100\times 100$, and $5$ pixels otherwise. The rest of the parameters were chosen according to the guidelines mentioned by Mukherjee and Qiu (2015) and Qiu and Yandell (1997).
@@ 3 Statistical Properties
Theorem 1: Assume that f has continuous first-order derivatives over $(0,1)\times(0,1)$ except on JLCs and the first order right and left derivative exists at the JLCs, $\epsilon_{ij}$ are independent and have distribution $N(0,\sigma^{2})$ with $0<\sigma^{2}<\infty$. If we have ${h}^{\ast}_{n}=O(n^{\beta})$ with $\beta<1$, $\alpha_{n}=O(1/\log\log n)$, $h_{n}=o(1)$, $1/nh_{n}=o(1)$ and the clustering bandwidth $\gamma_{n}<h_{n}\text{ with }\gamma_{n}=O(n^{-1}\log n)$, where $h^{\ast}_{n}$ is the bandwidth parameter for edge detection and $h_{n}$ is the bandwidth parameter for smoothing. Then, for any non-singular point $(x,y)$ with the property $d((x,y),(x^{*},y^{*}))>\epsilon$ for all singular points $(x^{*},y^{*})$, we have $\widehat{f}(x,y)=f(x,y)+O(\max(h_{n},\gamma_{n}^{1/2}))$. where $\widehat{f}(x,y)$ is the estimate of $f(x,y)$ produced by the proposed integrated method.
Sketches of the proof of the above theorem are based on Mukherjee and Qiu (2015) and Qiu and Yandell (1997). We provide the details in Appendix.
@@ 4 Numerical Studies
In this section, we present some numerical studies where we study the performance of the proposed integrated method which we call NEW in comparison with four competing methods. We consider three noise levels representing low, medium, and high, and various state-of-the-art denoising techniques for performance comparison. We use integrated root mean square error (IRMSE) to measure the performance of various image denoising methods which is defined as:
@@
In our numerical studies, we use estimated IRMSE by performing multiple instances of repeated independent simulations, and calculating the sample mean of root mean square error (RMSE) values.
@@ 4.1 A brief description of the competing methods
(ii) Steering kernel algorithm : The steering kernel method proposed by Takeda et al. (2007) uses an adaptive approach to denoise a given image. Instead of calculating the edge pixels directly it uses gradient information to calculates a direction which should be parallel to the nearby edge, and determines an elliptical neighborhood elongated in that direction. Since we incorporate a similar approach near the jump location curves of the given image, it is natural to include this method in our comparisons. The reader should note that the steering kernel method is an iterative method, i.e it applies the denoising method several times until the estimated mean squared error is small enough, whereas our proposed method is not iterative, hence computationally less expensive. We select the parameter of the steering kernel method as suggested by Takeda et al. (2007). We use the computer codes provided by Takeda (2007) for implementation.
@@
(iii) Jump preserving local linear kernel (JPLLK) method: This procedure proposed by Qiu (2009) involves a two-step process for surface reconstruction using local linear kernel smoothing. At a given point, a plane is fitted using local linear kernel smoothing to find a standard local linear kernel estimator of the image intensity function. The surrounding area is split into two sections by a line that passes through the point and is perpendicular to the gradient of the fitted plane. In each section, a half-plane is fitted using local linear kernel smoothing, resulting in two one-sided estimators of the image intensity function at that point. Subsequently, the best estimate among two one-sided estimates and the standard two-sided estimate is chosen as the final estimate at that point. This process is iterated twice to further smooth the surface. We select the bandwidth parameter of this smoothing procedure to get numerically optimal performance. Kang and Qiu (2024) provide computer codes for execution.
@@
(iv) Method by total variation regularization (TV): This is a very popular method used extensively in various applications. In this method, the total variation of an image is minimized using certain constraints related to the statistics of the noise. Rudin et al. (1992) show that reducing the total variation of an image removes noises while preserving essential structural details. Moreover, Chambolle and Pock (2011) develop a first order primal dual algorithm based on this method that works well for image denoising purposes alongside various other applications. We select parameters following the suggestion by Chambolle and Pock (2011). We use the computer codes provided by You (2021).
@@
It is to be noted that we do not consider any deep learning based denoising algorithms (e.g., Zamir et al. (2022)) as the competitor of the proposed method because they require a large set of images for training or learning. Hence, such type of denoising methods cannot be easily integrated with image monitoring procedures which itself is a computationally heavy problem. Moreover, the problem setup that we consider does not involve any training or learning on a large class of images. Therefore, comparing the proposed method with deep learning based denoising algorithms is rather unfair.
@@ 4.2 Simulation studies
We first construct a simple simulated image consisting of a square and a part of a circle. This image has various types of edges and edge structures such as linear edge, corner, intersection of a line with a curve and so forth. A noisy version of the simulated image is presented in the $(1,1)$-th panel of Figure 3. We choose such an image demonstrate the performance of the proposed integrated method in comparison with its competitors. The outcomes are presented in Table 1 and Figure 3.
@@
Proposed Algorithm Competing Methods
Resolution Noise SD NEW Steering Kernel Clustering Method JPLLK TV
@@
Figure 3: Denoising performance on simulated image with added Gaussian noise of 20 standard deviation. Images for top-left to bottom-right are the noisy image, denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK and TV, respectively.
Table 1 shows the performance of the concerned methods in terms of IRMSE. We run each of the concerned methods on images with different instances of noise $30$ times independently, and calculated the sample mean of RMSE as our estimated IRMSE. The numbers in parentheses represent the sample standard deviation of our estimates. Table 1 reflects the same phenomenon we observe from Figure 3. The numerical performance of the proposed method is quite similar to the clustering method, but much better than other competitors. It is expected to be similar to the clustering method as the proposed method is an improvement in terms of computational complexity. As seen from Table 1 and Figure 3, the proposed integrated method method performs slightly better in the regions where the true image is smooth which result in small improvement. In comparison with the steering kernel algorithm, the proposed method is better both visually and with respect to the estimated IRMSE values. JPLLK performs quite well in terms of smoothing visually, but it performs poorly where two jump location curves intersect which is reflected in the estimated IRMSE values. While the TV method performs well in low noise levels, it performs poorly in high noise levels. Our proposed integrated method performs better in comparison to its competitors especially when the noise level is high.
@@ 4.3 Performance comparison on real images
In this subsection, we consider two real images, one of a cityscape and another of a house to evaluate the performance of the denoising methods. Noisy versions of the the images are presented in the first panel of Figures 4 and 5. The outcomes are presented in Table 2, and Figures 4 and 5.
Proposed Algorithm Competing Methods
Image Noise SD NEW Steering Kernel Clustering Method JPLLK TV
@@
Figure 4: Performance comparison on the cityscape image with added Gaussian noise of 10 standard deviation. Images for top-left to bottom-right are the noisy image, denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK and TV, respectively.
Figure 5: Performance comparison on the house image with Gaussian added noise of 10 standard deviation. Images for top-left to bottom-right are the noisy image, denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK and TV, respectively.
From Table 2, we see the same patterns as in the case of the simulated image. The proposed algorithm outperforms the steering kernel method, JPLLK and TV in terms of estimated IRMSE for both of the images across all three noise levels that we consider. The performance of the proposed algorithm and the local clustering method continue to be similar both visually and in terms of estimated IRMSE. However, in several cases, the estimated IRMSE of the proposed method is smaller than that of the local clustering method. The steering kernel algorithm and TV method perform reasonably well in low noise level. The proposed method outperforms them in all cases with bigger margins at higher noise levels. In contrast, the performance of JPPLK is good at higher noise level but it perform rather poorly at lower noise level. Therefore, the proposed integrated algorithm performs well in nearly all situations without compromising the smoothing performance in the continuity regions while being less computationally expensive than the local clustering method.
Another important aspect to note is that the proposed algorithm performs better when the underlying edge structures of images are defined more clearly. This is because it relies heavily upon the edge detection algorithm we use. Therefore, in cases where edge pixels are not detected well, the proposed method method ends up smoothing those edges. However, when the noise level is very high, this issue is usually overshadowed due to heavy noise in comparison with the jump size.
@@ 5 Concluding Remarks
In this paper, we propose an integrated image denoising procedure which detects the edge pixels first, and then proceed to denoise using either adaptive kernel regression or local clustering depending on local information regarding edges. This paper shows better performance when compared to similar methods already in the literature. We also provide a theoretical justification that the image intensity function of the denoised image converges point-wise to the true image intensity function for large image size, under mild regularity assumptions. However, there are scopes for extending the proposed approach. In this paper, we consider the local clustering algorithm developed by Mukherjee and Qiu (2015) as baseline for denoising around edges. This can be improved by using more recent image denoising methods. Though it is restricted by the framework of the problem that we consider in this paper, one may expand this approach to use other types of denoising methods such as wavelet based methods, diffusion based methods, and many more. Another direction of potential improvement is the number of methods we can stitch together using this approach. In this paper, we divide the problem in just two cases, namely regions near edges and regions far away from edges. However, in future, one may sub-divide the problem into more parts, and thus have an assortment of methods that perform well in specific scenarios which all together to produce superior results. One immediate use of the proposed image denoising approach is in developing methods for image comparisons and image monitoring.
@@ Appendix A Appendix
$S_{\epsilon}$ =$=\left\{(x,y):(x,y)\in\Omega,d_{E}\left((x,y),\left(x^{*},y^{*}\right)\right)\leq\epsilon\text{ for a singular point }\left(x^{*},y^{*}\right)\in D\right\},$
$\Omega_{\bar{S},\epsilon}$ $=\Omega_{\epsilon}\backslash S_{\epsilon}.$
@@
Lemma A.1
Assume that $f(x,y)$ has continuous first order partial derivatives over $(0,1)\times(0,1)$ except on the JLCs and it has the first order right and left partial derivatives at the JLCs. And also $\lim_{n\rightarrow\infty}h^{\ast}=\infty$ and $\lim_{n\rightarrow\infty}h^{\ast}/n=0$ where $h^{\ast}$ is the window width. If there is no jump in $N\left(x_{i},y_{j}\right)$, then
@@
[…]
where $\left(\tilde{x}_{i},\tilde{y}_{j}\right)$ is some point around $\left(x_{i},y_{j}\right)$ on the same side of the JLC as $\left(x_{i},y_{j}\right)$ and the distance between $\left(\tilde{x}_{i},\tilde{y}_{j}\right)$ and $\left(x_{i},y_{j}\right)$ tends to zero; $C(i,j),C_{x}(i,j)$, and $C_{y}(i,j)$ are the jump magnitudes of $f(x,y)$ and its first order $x$ and y partial derivatives; $h_{1}^{(i,j)}$ and $h_{2}^{(i,j)}$ are two constants satisfying
@@
$\gamma_{1}$ and $\gamma_{2}$ are two constants between $-1$ and 1.
Lemma A.2
Besides the conditions stated above, if there are only finite number of singular points on JLCs and $\alpha_{n}$ is chosen such that
  1. i.
@@
Lemma A.3
Assume that $f$ has continuous first-order derivatives over $(0,1)\times(0,1)$ except on the JLCs, its first order derivatives have one-sided limits at non-singular points of the JLCs, $\left\{\varepsilon_{ij}\right\}$ are i.i.d. and have the common distribution $N\left(0,\sigma^{2}\right)$, where $0<\sigma<\infty$, $h_{n}=o(1),1/nh_{n}=o(1),u_{n}=\kappa+\delta$, where $\kappa=\left(\frac{\phi^{2}(0)}{\Phi(0)(1-\Phi(0))}\right)/\left[1-\left(\frac{\phi^{2}(0)}{\Phi(0)(1-\Phi(0))}\right)\right],\delta$ is any positive number, and $\phi$ and $\Phi$ are the probability density function and the cumulative distribution function of the $N(0,1)$ distribution, respectively. If we further assume that $B_{n}=$ $O\left(h_{n}^{1/2}\right)$, then for any non-singular $(x,y)\in\Omega_{\bar{S},\epsilon}$, we have $\widehat{f}(x,y)=f(x,y)+O\left(h_{n}^{1/2}\right)$, a.s.
@@
From Lemmas 1 and 2, we can say that if we choose $\alpha_{n}=O(1/log(log(n))$ and $h^{\ast}=O(n^{\beta})$ where $\beta<1$, all the above conditions are satisfied for edge detection.
Now suppose we are using $h^{\ast}_{n}$ window width and $\alpha_{n}$ confidence level for edge detection and $h_{n}$ window width with $h_{n}=o(1)$ and $\gamma_{n}$ as bandwidth parameter for clustering according to Lemma 3 for the smoothing process, all of the conditions required regarding the parameters of above lemmas are satisfied. We now proceed to prove Theorem 1.
@@
    Let $J$ denote the points on JLCs and $\hat{J_{n}}$ denote the detected edge pixels. Suppose we are given the point $(\tilde{x},\tilde{y})$.
    Suppose the point is not on $J$. Then as $\mathcal{H}(J,\hat{J}_{n})\rightarrow 0$, for large enough n it is not in $\hat{J}_{n}$ with probability 1. So without loss of generality we can assume it is not on $\hat{J}_{n}$ Suppose the nearest point on $\hat{J}_{n}$ is $(x^{\ast},y^{\ast})$.
    There exists $N$ such that for all $n>N$, we have $\mathcal{H}(J,\hat{J_{n}})<\frac{\gamma_{n}}{2}$. So we can say $\forall n>N$ there is a circle around $(x^{\ast},y^{\ast})$ of radius $d((x^{\ast},y^{\ast}),(\tilde{x},\tilde{y}))-\gamma_{n}$ which does not touch $\hat{J_{n}}$.
    Now, for choosing the second closest point in our algorithm (point $P_{2}$ in Figure 2), we are doing the same as above while restricting the image in that strip. So we can take $\hat{\Omega}=\Omega\cap S$ and $\hat{J}^{{}^{\prime}}_{n}=\hat{J}_{n}\cap S$, where $S$ is the strip chosen using the first step. So from the fact that the ellipse we define is completely inside the strip and using the same argument as choosing the first point we can say that the ellipse we are choosing has no point which are on JLC. So. $f(x,y)$ has continuous first order derivatives inside the elliptical neighborhood $E$. As the maximum possible axis length, $h_{n}\rightarrow 0$, the Taylor series converges, Hence $\hat{f}(\tilde{x},\tilde{y})\rightarrow f(\tilde{x},\tilde{y})$ i.e $\widehat{f}(\tilde{x},\tilde{y})=f(\tilde{x},\tilde{y})+O\left(h_{n}-\gamma_{n}\right)=f(\tilde{x},\tilde{y})+O\left(h_{n}\right)$, a.s.
    Now if $(\tilde{x},\tilde{y})$ is on J. Then as $\mathcal{H}(J,\hat{J}_{n})\rightarrow 0$, we can say, $d((\tilde{x},\tilde{y}),\hat{J}_{n})<\gamma_{n}$ for large enough n. So, for large enough n we will be using the method proposed by Mukherjee and Qiu (2015). So, by Lemma 3, we get, $\widehat{f}(\tilde{x},\tilde{y})=f(\tilde{x},\tilde{y})+O\left(\gamma_{n}^{1/2}\right)$ a.s. Hence Theorem 1 is proved. $\blacksquare$

### v3

@@ 2.3 Selection of procedure parameters
For our proposed algorithm, we need four parameters: the bandwidth and cutoff for the edge detection step, and bandwidth parameters for the local clustering step and smoothing step, and one tuning parameter for the smoothing step after local clustering. In our implementation, we introduce only one new parameter, the maximum possible axis length of the elliptical neighborhood. From numerical experience, we set this parameter to $6$ pixels for images having resolution less than $100\times 100$, and $10$ pixels for images with higher resolution. Similarly, the bandwidth for the clustering method is chosen to be the same as the minimum distance $\gamma$ to apply the method, which we choose to be $3$ pixels for images with resolution less than $100\times 100$, and $5$ pixels otherwise. The rest of the parameters are chosen according to the guidelines mentioned by Mukherjee and Qiu (2015) and Qiu and Yandell (1997).
2.4 Computational efficiency
As our algorithm integrates two different algorithms with distinct computation times, comparison of their computational efficiencies is essential. Suppose, $\mathcal{C}_{1}$ denotes the computation time of the edge detection algorithm, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ denote the computation times of the regression method and the clustering method per pixel, respectively. Hence, the total computation time $\mathcal{C}_{new}$ of the integrated algorithm is given by:
$\mathcal{C}_{new}=\mathcal{C}_{1}+\sum\limits_{x,y}(P(x,y)\mathcal{C}_{2}+(1-P(x,y))\mathcal{C}_{3}),$
where $P(x,y)$ represents the probability that the nearest edge pixel to the point $(x,y)$ lies at a distance greater than $\gamma$. Noting that $\mathcal{C}_{1}+n^{2}\min(\mathcal{C}_{2},\mathcal{C}_{3})\leq\mathcal{C}_{new}\leq\mathcal{C}_{1}+n^{2}\max(\mathcal{C}_{2},\mathcal{C}_{3})$, we conclude that the computation time of the proposed integrated method lies between those of a single iteration of the two baseline algorithms. Moreover, the regression based smoothing by Takeda et al. (2007) involves multiple iterations, hence its computation time is $\mathcal{T}n^{2}\mathcal{C}_{2}$, $\mathcal{T}$ being the number of iterations. The computation time of the local clustering method by Mukherjee and Qiu (2015) is $\left(\mathcal{C}_{1}+n^{2}\mathcal{C}_{3}\right)$ with $\mathcal{C}_{3}$ being slightly larger than $\mathcal{C}_{2}$, but of similar order with respect to $n$. Hence, for sufficiently large $\mathcal{T}$ and $n$, $\mathcal{C}_{new}$ is smaller than both $\mathcal{T}n^{2}\mathcal{C}_{2}$ and $\left(\mathcal{C}_{1}+n^{2}\mathcal{C}_{3}\right)$.
Although the proposed integrated algorithm involves a two-step procedure, apart from initial edge detection, the rest of the algorithm can be performed independently on each pixel. Therefore, the second step of this integrated algorithm can be parallelized. Furthermore, by choosing a computationally lightweight edge detection algorithm, we can significantly reduce the overall computation time.
@@ 3 Statistical Properties
Theorem 1: Assume that f has continuous first-order derivatives over $(0,1)\times(0,1)$ except on JLCs, the first order one-sided derivatives exist at the JLCs, $\epsilon_{ij}$ are independent and have distribution $N(0,\sigma^{2})$ with $0<\sigma^{2}<\infty$. If we have ${h}^{\ast}_{n}=O(n^{\beta})$ with $\beta<1$, $\alpha_{n}=O(1/\log\log n)$, $h_{n}=o(1)$, $1/nh_{n}=o(1)$, and the clustering bandwidth $\gamma_{n}<h_{n}\text{ with }\gamma_{n}=O(n^{-1}\log n)$, where $h^{\ast}_{n}$ is the bandwidth parameter for edge detection, and $h_{n}$ is the bandwidth parameter for smoothing. Then, for any non-singular point $(x,y)$ with the property $d((x,y),(x^{*},y^{*}))>\epsilon$ for all singular points $(x^{*},y^{*})$, we have $\widehat{f}(x,y)=f(x,y)+O(\max(h_{n},\gamma_{n}^{1/2}))$. where $\widehat{f}(x,y)$ is the estimate of $f(x,y)$ produced by the proposed integrated method.
Sketches of the proof of the above theorem are based on Mukherjee and Qiu (2015) and Qiu and Yandell (1997). We provide the details in the Appendix.
Remark: Although we make normality assumption of the additive noise for simplicity of the proof, it is not critically important. As long as the density of the error distribution is symmetric around zero, and has finite moments at least up to a certain order, point-wise convergence of $\widehat{f}(x,y)$ holds. If these assumptions on the error distribution do not hold, then we need to use other techniques, such as noise-adaptive transforms or pre-processing for non-Gaussian noise.
@@ 4 Numerical Studies
In this section, we present several numerical experiments where we study the performance of the proposed integrated method which we call NEW, in comparison with four competing methods. We consider three noise levels representing low, medium, and high, and various state-of-the-art denoising techniques for performance comparison. We use integrated root mean square error (IRMSE) to measure the performance of various image denoising methods which is defined as:
@@
In our numerical studies, we estimate IRMSE by performing multiple instances of repeated independent simulations, and calculating the sample mean of root mean square error (RMSE) values:
@@ 4.1 A brief description of the competing methods
(ii) Steering kernel algorithm: The steering kernel method proposed by Takeda et al. (2007) uses an adaptive approach to denoise a given image. Instead of calculating the edge pixels directly, it uses gradient information to find a direction which should be parallel to the nearby edge, and determines an elliptical neighborhood elongated in that direction. Since we incorporate a similar approach near the jump location curves of the given image, it is natural to include this method in our comparison study. It is to be noted that the steering kernel method is an iterative method, i.e., it applies the denoising method several times until the estimated mean squared error is small enough, whereas our proposed method is not iterative, and hence computationally less expensive. We select the parameters of the steering kernel method as suggested by Takeda et al. (2007). We use the computer codes provided by Takeda (2007) for implementation.
@@
(iii) Jump preserving local linear kernel (JPLLK) method: This procedure proposed by Qiu (2009) involves a two-step process for surface reconstruction using local linear kernel smoothing. At a given point, a plane is fitted using local linear kernel smoothing to find a standard local linear kernel estimator of the image intensity function. The surrounding area is split into two sections by a line that passes through the point and is perpendicular to the gradient of the fitted plane. In each section, a half-plane is fitted using local linear kernel smoothing, resulting in two one-sided estimators of the image intensity function at that point. Subsequently, the best estimate among the two one-sided estimates, and the standard two-sided estimate is chosen as the final estimate at that point. This process is iterated twice to further smooth the surface. We select the bandwidth parameter of this smoothing procedure to get numerically optimal performance. Kang and Qiu (2024) provide computer codes for execution.
@@
(iv) Method by total variation regularization (TV): This is a very popular method used extensively in various applications. In this method, the total variation of an image is minimized using certain constraints related to the statistics of the noise. Rudin et al. (1992) show that reducing the total variation of an image removes noise while preserving essential structural details. Moreover, Chambolle and Pock (2011) develop a first order primal dual algorithm based on this method that works well for image denoising purposes alongside various other applications. We select parameters following the suggestion by Chambolle and Pock (2011). We use the computer codes provided by You (2021).
@@
It is to be noted that we do not consider any deep learning based denoising algorithms (e.g., Zamir et al. (2022)) as the competitor of the proposed method because they require a large set of images for training or learning. Hence, such type of denoising methods cannot be easily integrated with image monitoring procedures which itself is a computationally heavy problem. Moreover, the problem setup that we consider does not involve any training or learning on a large class of images. Although the self-supervised learning methods Quan et al. (2020); Chen et al. (2024) simulate training data from the given noisy image, computational complexity still remains as an issue if we wish to use the denoising method in subsequent analysis such as image monitoring. Therefore, comparing the proposed method with deep learning based denoising algorithms is rather unfair.
@@ 4.2 Simulation studies
We first construct a simple simulated image consisting of a square and a part of a circle. This image has various types of edges and edge structures such as linear edge, corner, intersection of a line with a curve and so forth. A noisy version of the simulated image is presented in the $(1,1)$-th panel of Figure 3. We choose such an image to demonstrate the performance of the proposed integrated method in comparison with its competitors. The outcomes are presented in Table 1 and Figure 3.
@@
Competing Methods
Resolution Noise Proposed Steering Kernel Clustering JPLLK TV
@@
Figure 3: Denoising performance on the simulated image with additive Gaussian noise of 2020 standard deviation. Panels from the top-left to the bottom-right show the noisy image, and the denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK, and TV, respectively.
Table 1 shows the performance of the concerned methods in terms of IRMSE. We run each of the concerned methods on images with different instances of noise $30$ times independently, and calculate the sample mean of RMSE as our estimated IRMSE. The numbers in parentheses show the standard error of the estimates. Table 1 reflects the same phenomenon we observe from Figure 3. The numerical performance of the proposed method is quite similar to the clustering method, but much better than other competitors. It is expected to be similar to the clustering method as the proposed method is an improvement in terms of computational complexity. As seen from Table 1 and Figure 3, the proposed integrated method performs slightly better in the regions where the true image is smooth. This results in small improvement in the estimated IRMSE. In comparison with the steering kernel algorithm, the proposed method is better both visually and with respect to the estimated IRMSE values. JPLLK performs quite well in terms of smoothing visually, but it performs poorly where two jump location curves intersect which is reflected in the estimated IRMSE values. While the TV method performs well in low noise levels, it performs poorly in high noise levels. Our proposed integrated method performs better in comparison to its competitors especially when the noise level is high.
The performance of the proposed integrated method along with the baseline methods are very similar when the noise distribution is non-Gaussian but symmetric around zero with finite first few moments. Performances of the jump regression based image denoising methods have been explored in the literature in presence of such kinds of non-Gaussian noise, and these performances are very similar to the Gaussian noise cases Mukherjee and Qiu (2015); Qiu and Mukherjee (2010). These observations are in line with our remark after Theorem 1.
@@ 4.3 Performance comparison on real images
In this subsection, we consider two real images, one of a cityscape, and another of a house to evaluate the performance of the denoising methods. Noisy versions of the images are presented in the first panel of Figures 4 and 5. The outcomes are presented in Table 2, and Figures 4 and 5.
Competing Methods
Image Noise Proposed Steering Kernel Clustering JPLLK TV
@@
Figure 4: Performance comparison on the cityscape image with additive Gaussian noise of 10 standard deviation. Panels from the top-left to the bottom-right show the noisy image, and the denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK, and TV, respectively.
Figure 5: Performance comparison on the house image with additive Gaussian added noise of 10 standard deviation. Panels from the top-left to the bottom-right show the noisy image, and the denoised images by the proposed method, clustering algorithm, steering kernel, JPLLK, and TV, respectively.
From Table 2, we observe similar patterns as in the case of the simulated image. The proposed algorithm outperforms the steering kernel method, JPLLK, and TV in terms of estimated IRMSE for both of the images across all three noise levels that we consider. The performance of the proposed algorithm and the local clustering method continue to be similar both visually and in terms of estimated IRMSE. In fact, the estimated IRMSE of the proposed method is smaller than that of the local clustering method in several cases. The steering kernel algorithm and TV method perform reasonably well when the noise level is low. The proposed method outperforms them in all cases with bigger margins at higher noise levels. In contrast, the performance of JPPLK is good at higher noise level but it perform rather poorly at lower noise level. Therefore, the proposed integrated algorithm performs well in nearly all situations without compromising the smoothing performance in the continuity regions while being less computationally expensive than the local clustering method.
Another important aspect to note is that the proposed algorithm performs better when the underlying edge structures of images are defined more clearly. This is because it relies heavily upon the edge detection algorithm we use. Therefore, in cases where edge pixels are not detected well, the proposed method ends up smoothing those edges. However, when the noise level is very high, this issue is usually overshadowed due to heavy noise in comparison with the jump size.
@@ 5 Concluding Remarks
In this paper, we propose an integrated image denoising procedure which detects the edge pixels first, and then proceed to denoise using either adaptive kernel regression or local clustering depending on local information regarding edges. This paper shows better performance when compared to similar methods already in the literature. We also provide a theoretical justification that the image intensity function of the denoised image converges point-wise to the true image intensity function for large image size, under mild regularity assumptions. However, there are scopes for extending the proposed approach. In this paper, we consider the local clustering algorithm developed by Mukherjee and Qiu (2015) as baseline for denoising around edges. This can be improved by using more recent image denoising methods. Though it is restricted by the framework of the problem that we consider in this paper, one may expand this approach to use other types of denoising methods such as wavelet based methods, diffusion based methods, and many more. Another direction of potential improvement is the number of methods we can stitch together using this approach. In this paper, we divide the problem into just two parts, namely smoothing near the edges, and smoothing far away from the edges. However, in future research, one may divide the problem into more parts, and thus have an assortment of methods that perform well in specific scenarios which altogether produce superior overall result. One immediate application of the proposed image denoising method is in designing techniques for image comparison and image monitoring.
@@ Appendix A Appendix
$S_{\epsilon}$ $=\left\{(x,y):(x,y)\in\Omega,d_{E}\left((x,y),\left(x^{*},y^{*}\right)\right)\leq\epsilon\text{ for a singular point }\left(x^{*},y^{*}\right)\in D\right\},$
[…]
@@
Lemma 1: Assume that $f(x,y)$ has continuous first order partial derivatives over $(0,1)\times(0,1)$ except on the JLCs, and its first order one-sided partial derivatives at the JLCs exist. Assume that $\lim_{n\rightarrow\infty}h^{\ast}=\infty$ and $\lim_{n\rightarrow\infty}h^{\ast}/n=0$, where $h^{\ast}$ is the window width. If there is no jump in $N\left(x_{i},y_{j}\right)$, then
@@
$\hat{\beta}_{2}^{(i,j)}=f_{y}^{\prime}\left(\tilde{x}_{i},\tilde{y}_{j}\right)+h_{2}^{(i,j)}C(i,j)+\Gamma_{2}C_{y}(i,j)+O\left(\frac{n\sqrt{\log\log h^{\ast}}}{{h^{\ast}}^{2}}\right)\quad\text{a.s.,}$
where $\left(\tilde{x}_{i},\tilde{y}_{j}\right)$ is some point around $\left(x_{i},y_{j}\right)$ on the same side of the JLC as $\left(x_{i},y_{j}\right)$, and the distance between $\left(\tilde{x}_{i},\tilde{y}_{j}\right)$ and $\left(x_{i},y_{j}\right)$ converges to zero; $C(i,j),C_{x}(i,j)$, and $C_{y}(i,j)$ are the jump magnitudes of $f(x,y)$ and its first order $x$ and y partial derivatives; $h_{1}^{(i,j)}$ and $h_{2}^{(i,j)}$ are two constants satisfying
@@
$\Gamma_{1}$ and $\Gamma_{2}$ are two constants between 1-1 and +1+1.
Lemma 2: Besides the conditions stated above, if there are only finite number of singular points on JLCs, and $\alpha_{n}$ is chosen such that
@@
Lemma 3: Assume that $f$ has continuous first-order derivatives over $(0,1)\times(0,1)$ except on the JLCs, its first order derivatives have one-sided limits at non-singular points of the JLCs, $\left\{\varepsilon_{ij}\right\}$ are i.i.d. and have the common distribution $N\left(0,\sigma^{2}\right)$, where $0<\sigma<\infty$, $h_{n}=o(1)$, $1/nh_{n}=o(1)$, $u_{n}=\kappa+\delta$, where $\kappa=\left(\frac{\phi^{2}(0)}{\Phi(0)(1-\Phi(0))}\right)/\left[1-\left(\frac{\phi^{2}(0)}{\Phi(0)(1-\Phi(0))}\right)\right]$, $\delta$ is any positive number; $\phi$ and $\Phi$ are the probability density function and the cumulative distribution function of the $N(0,1)$ distribution, respectively. If we further assume that $B_{n}=$ $O\left(h_{n}^{1/2}\right)$, then for any non-singular $(x,y)\in\Omega_{\bar{S},\epsilon}$, we have $\widehat{f}(x,y)=f(x,y)+O\left(h_{n}^{1/2}\right)$, a.s.
@@
From Lemmas 1 and 2, we can say that if we choose $\alpha_{n}=O(1/log(log(n))$ and $h^{\ast}=O(n^{\beta})$ where $\beta<1$, all of the above conditions are satisfied for edge detection.
Suppose that we select the bandwidth $h^{\ast}_{n}$ and the significance level $\alpha_{n}$ for edge detection; the bandwidths $h_{n}$ and $\gamma_{n}$ for smoothing and clustering, respectively, and the tuning parameter $B_{n}$ for smoothing, according to Lemma 3. Now that all required conditions for the parameters of above lemmas are satisfied, we now proceed to prove Theorem 1.
@@
[…]
    Suppose that the point is not on $J$. Then, as $\mathcal{H}(J,\hat{J}_{n})\rightarrow 0$, for large $n$, it is not in $\hat{J}_{n}$ with probability 1. Therefore, without any loss of generality, we can assume that it is not on $\hat{J}_{n}$. Suppose that its nearest point on $\hat{J}_{n}$ is $(x^{\ast},y^{\ast})$.
[…]
    Theorem 1 is thus proved.